OpenStudy (anonymous):

prove arcsinh 3/4 + arcsinh 5/12 =arcsinh 4/3

zepdrix (zepdrix):

Hey there :) Hmm this problem looks kinda fun.

OpenStudy (anonymous):

hello, its fun for me once i get the answer

zepdrix (zepdrix):

Are you familiar with this definition of arcsinh?\[\large\rm arcsinh(x)=\ln\left(x+\sqrt{x^2+1}\right)\]If not, we can derive that equation from sinh(x) and take inverse. Otherwise we can start from this point.

OpenStudy (anonymous):

yeah, i do learnt about it

zepdrix (zepdrix):

\[\rm arcsinh\left(\frac{3}{4}\right)+arcsinh\left(\frac{5}{12}\right)\]So converting to log stuff,\[\rm \ln\left(\frac{3}{4}+\sqrt{\left(\frac{3}{4}\right)^2+1}\right)+\ln\left(\frac{5}{12}+\sqrt{\left(\frac{5}{12}\right)^2+1}\right)\]First step is simply getting everything set up correctly. Ok with that? :o

OpenStudy (anonymous):

yup

zepdrix (zepdrix):

From there it should be easy peasy :) Just a bunch of algebra (leave everything as fractions, no decimals), then one application of log rule: \(\rm log(a)+log(b)=log(a\cdot b)\)

zepdrix (zepdrix):

Square your things,\[\rm \ln\left(\frac{3}{4}+\sqrt{\frac{9}{16}+1}\right)+\ln\left(\frac{5}{12}+\sqrt{\frac{25}{144}+1}\right)\]and then get common denominators, ya? :)

OpenStudy (anonymous):

just a question, when should we used the triangle and when we should apply the ln form

zepdrix (zepdrix):

Ooo I don't really remember triangle method for hyperbolic, I'm not sure I would be able to answer that :(

OpenStudy (anonymous):

alright, thanks for helping :)

zepdrix (zepdrix):

c: